OpenStudy (jessiegonzales):

CHECK MY ANSWER? MEDAL!!! For questions 11 and 12, find the volume of the cylinder in terms of pi.

OpenStudy (jessiegonzales):

52.8m^3 58.08pi m^3 127.78 pi m^3<---- 29.04pi m^3

sammixboo (sammixboo):

Do you know the formula of the volume of a cylinder?

OpenStudy (jessiegonzales):

v=pi r^2h

sammixboo (sammixboo):

It is \(\rm \large V~=~\pi~r^2h\), correct. We know that the radius (r) is 2.2 and the height (h) is 12, given in the picture.. Plug in the values into the formula first: \(\rm \large V~=~\pi~2.2^2\times h\) So now we know \(\rm \large V~=~\pi~2.2^2\times h\) First we must do the exponents, so what is \(\rm 2.2^2\)?

sammixboo (sammixboo):

Agh sorry. I didn't fill in the formula all he way.. It is \(\rm \large V~=~\pi~2.2^2\times 12\) Still do the exponent first..

OpenStudy (jessiegonzales):

4.84

sammixboo (sammixboo):

Correct! So plug in 4.84 where \(\rm 2/2^2\) was in the equation... So now we have \(\rm \large V~=~\pi~4.84\times 12\) Next to the multiplication, so what is 4.84 x 12?

sammixboo (sammixboo):

2.2^2* I am soooo sorry! Today is Typo Day for me <_>

OpenStudy (jessiegonzales):

58.08

sammixboo (sammixboo):

Correct, so now we have \(\rm \large V~=~\pi~58.08\), or \(\rm \large V~=~58.08\pi\)

sammixboo (sammixboo):

So that would be our answer

OpenStudy (jessiegonzales):

mind helping with another one

sammixboo (sammixboo):

Yeah, sure :)

sammixboo (sammixboo):

Don't mind my typos, btw ;(

OpenStudy (jessiegonzales):

h=6in r=3in

sammixboo (sammixboo):

So you are trying to find the volume?

OpenStudy (jessiegonzales):

yes

sammixboo (sammixboo):

Okay o get the volume, V = pi r^2*h And plug in the information

sammixboo (sammixboo):

All we know is the r and the h, so let's plug those in... V = pi 3^2*6 First do the exponent, 3^2, so what is 3^2?

OpenStudy (jessiegonzales):

54pi in^2 is the answer

sammixboo (sammixboo):

Yeppers, because 3^2 = 9 and 9*6 = 54, so all we have left is 54pi